OpenStudy (anonymous):

cos(180-x)=-cos x verify it with identities. @StudyGurl14 can you possibly help me out? i will give medal to best answer! thanks

OpenStudy (dumbcow):



OpenStudy (anonymous):

it is a verification problem, identities need to be used to make one side identical to the other. the picture does not verify it.

OpenStudy (dumbcow):

ok but the picture does indeed show that they are equal using the unit circle anyway you need to use this property for difference of angles \[\cos(a-b) = \cos a \cos b + \sin a \sin b\]

OpenStudy (anonymous):

ohhhh thanks! i completely forgot about that identity!